OpenStudy (anonymous):

in a recent year the total sales at restaurant A were 400,000 more than restaurant B the total were 625,000,000. what were the sales for each?

OpenStudy (anonymous):

A=B+400,000 B+A=626,000,000 B+B+400,000=626,000,000 solve for B and Then sub B into A=B+400,000 to find A

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

625,000,000..... just a little typo :)

OpenStudy (amistre64):

or 5 million little typos :)

OpenStudy (jeremy):

b+400000+b=625000000 2b=624600000 b is =312300000 and a is=312700000